Particle moves up in air with initial velocity 100 m/s makes angle 60° with horizontal. The radius of curvature at t = 1 s is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
At t=1 s,
Vx = 100cos60 = 50 m/s
and
Vy = 100sin60 – g*1 = > 50root(3) – 10 = 75 m/s.
Vnet = root(50^2 + 75^2) => 90.13 m/s.
Radii of curvature would be = Perpendicular velocity to g/g => Vp/g
Therefore, Vp = 90.13cos60 = > 45.06 approx.
 
Hence,
ROC = 45.06/10 = > 4.50 approx.